Recent publications of Endre Csáki

Csáki, E., Csörgő, M. and Kulik, R.: Strong approximations for long memory sequences based partial sums, counting and their Vervaat processes. Under review

Csáki, E. and Csörgő, M.: On Bahadur-Kiefer type processes for sums and renewals in dependent cases. In: Mathematical Statistics and Limit Theorems: Conference in Honor of Paul Deheuvels. Springer, New York, to appear.

Csáki, E., Földes, A. and Révész, P.: Some results and problems for anisotropic random walk on the plane. In: Asymptotic Laws and Methods in Stochastics: A Volume in Honour of Miklós Csörgő. Fields Institute Communications, Springer, New York, to appear.

Csáki, E., Csörgő, M. Földes, A. and Révész, P.: Strong limit theorems for anisotropic random walks on Z^{2}. Periodica Math. Hungar. 67 (2013) 71–94.

Csáki, E., Csörgő, M. Földes, A. and Révész, P.: Random walk on half-plane half-comb structure. Ann. Mathematicae et Informaticae 39 (2012), 29–44.

Csáki, E., Csörgő, M. Földes, A. and Révész, P.: On the local time of random walk on the 2-dimensional comb. Stoch. Process. Appl. 121 (2011), 1290–1314.

Csáki, E., Csörgő, M. and Kulik, R.: On Vervaat processes for sums and renewals in weakly dependent cases. Dependence in Probability, Analysis and Number Theory. A Volume in Memory of Walter Philipp, edited by I. Berkes, R.C. Bradley, H. Dehling, M. Peligrad, R. Tichy. Kendrick Press, 2010, pp. 145–156.

Csáki, E., Csörgő, M., Földes, A. and Révész, P.: On the supremum of iterated local time. Publ. Math. Debrecen 76 (2010), 255–270.

Csáki, E., Földes, A. and Révész, P.: On the number of cutpoints of the transient nearest neighbor random walk on the line. J. Theoret. Probab. 23 (2010), 624– 638.